The slope of canal for a discharge of 300 cumecs should be

a.1in 4000

b.1in 6000

c.1in 8000

d.1in 10000
Answer #1

For a discharge of 300 m​​​​​​3​​​/s, slope of canal bed should be c) 1 in 8000.

Gerald Lacey was a chief engineer of UP irrigation department in INDIA. He proposed a theory in 1919, popularly known as 'Lacey's Regime Theory'. He gave certain regime relations based on his findings for unlined alluvial channels.

Among those relations, equation for bed slope (S) is given by:

S = \frac{f^{5/3}}{3340Q^{1/6}}

where,

f = silt factor = usually taken as 1, when sufficient data (average particle size is not available).

Q = Discharge in cumec

So,

For f = 1 and Q = 300 cumec,

S = \frac{1^{5/3}}{3340X300^{1/6}}=0.00012 = \frac{1}{8333} \sim \frac{1}{8000} approx